# DMesh++: An Efficient Differentiable Mesh for Complex Shapes

### v1

@@ 6 Discussion
In this work, we presented DMesh++, a novel approach to handle mesh in a differentiable manner.
As a core principle, we first introduced the Minimum-Ball algorithm (Sec. 3.2). This algorithm allows us to increase the mesh complexity in a computationally efficient manner, enabling the handling of complex shapes. Specifically, we demonstrated that this algorithm has a time complexity of $O(\log N)$, where $N$ is the number of points. This efficiency is achieved because the Minimum-Ball algorithm considers only the local point configuration near a face to determine the face’s existence. In contrast, DMesh [37] evaluates the global point configuration, resulting in a time complexity of $O(N)$, which is significantly more costly as $N$ grows.
Under the Minimum-Ball condition, we select faces from the Delaunay Triangulation (DT) that satisfy this criterion. Among these, we identify faces whose vertices all have a real value ($\psi$) of 1. These faces constitute the “real” part and form the final mesh. The remaining faces in the DT, which do not satisfy this criterion, are labeled as the “imaginary” part and are excluded from the final mesh (Fig. 1).
Additionally, we presented the Reinforce-Ball algorithm (Sec. 4.2) for 2D reconstruction tasks. This algorithm allows us to decrease the mesh complexity, producing a mesh that adapts to local geometry. Like the Minimum-Ball algorithm, the Reinforce-Ball algorithm evaluates only the local neighborhood of a face to determine its existence.
To summarize, these two algorithms enable us to achieve both computational efficiency and mesh efficiency. In this appendix, we provide detailed explanations and further discussions on various aspects of DMesh++.
@@ 7.1 Algorithm
We formally describe the Minimum-Ball algorithm in Algorithm 1. Below, we provide a line-by-line explanation of the algorithm:
@@ 7.2 Minimum-Ball computation
$B_{F}^{c}\big{|}_{d=3}=p_{1}+\frac{||d_{2}||^{2}(d_{1}\times d_{2})\times d_{1}+||d_{1}||^{2}(d_{2}\times d_{1})\times d_{2}}{2||d_{1}\times d_{2}||^{2}},$ (11)
@@ 7.3 Sigmoid coefficient $\alpha_{min}$
As illustrated in Figure 4, the reconstruction process starts from a fixed triangular (2D) or tetrahedral (3D) grid (Fig. 13). At the initial state, every face in the grid satisfies the Minimum-Ball condition (Appendix 8.2). Notably, every interior face in the grid shares a common signed distance $d_{common}>0$. Let us denote $x$ as the edge length of the grid, applicable for both 2D and 3D cases. Then, the common signed distance can be computed as follows:
@@ 7.4 Nearest neighbor caching
During the subsequent optimization steps, for a given face $F$, we compute the distance from $B_{F}^{c}$ to the cached $K$-nearest neighbors in $\mathbb{P}$ and select the nearest neighbor from the cache to compute the signed distance $d(B_{F},\mathbb{P})$. This mechanism is described in detail in Algorithm 2 and Appendix 8.3 in the context of point position optimization. In our experiments for 3D multi-view reconstruction, we set $n_{0}=2000$, $n_{1}=50$, and $K=10$.
8 Details about Reconstruction Pipeline
In this section, we provide implementation details for each step in the optimization pipeline (Fig. 4). Before delving into these details, we revisit the reconstruction loss used in our framework.
8.1 Reconstruction Loss ($L_{recon}$)
@@
For rendering probabilistic faces, we interpret each face’s existence probability as its opacity, following [37]. We use a differentiable renderer of [37], but enhance it with anti-aliasing implemented in CUDA (Appendix 11). This renderer is built upon the efficient Gaussian Splatting renderer proposed by [14]. While it does not perform exact depth ordering between faces, it efficiently handles a large number of semi-transparent faces, making it significantly faster than renderers that perform exact depth ordering.
To compute the rendering loss, we render the mesh from multiple viewpoints and compare the rendered images to ground truth images using the $L_{1}$ loss.
8.2 Step 1: Real Value Initialization
@@
In this step, we initialize the point-wise real values ($\psi$). To achieve this, we first organize the points in a regular grid, ensuring that every face in the grid satisfies the Minimum-Ball condition (Definition 3.1). This guarantees that the faces observed in this step will also be observable in the subsequent step, where the Minimum-Ball algorithm determines face existence. For $d=2$, this condition is satisfied by forming every triangle in the grid as an equilateral triangle. For $d=3$, we use a body-centered cubic lattice [3]. The grids are illustrated in Fig. 13.
With these fixed points and faces, we determine which faces to include in the final mesh by minimizing the reconstruction loss $L_{recon}$. Specifically, for each face $F$ in the grid, we assign an optimizable variable $\Xi(F)\in[0,1]$, representing the face’s existence probability. We then optimize $\Xi(F)$ to minimize $L_{recon}$. To reduce the number of redundant faces, we add a small regularization term $L_{real}$:
$L_{real}=\frac{1}{|\mathbb{F}|}\sum_{F\in\mathbb{F}}\Xi(F),$ (15)
where $\mathbb{F}$ is the set of faces in the grid. The total loss optimized in this step is:
$L=L_{recon}+\alpha_{real}\cdot L_{real}.$ (16)
We set $\alpha_{real}=10^{-4}$.
After optimization, we collect faces with probabilities larger than $0.01$ to ensure that as many faces as possible are available for the next optimization step, thereby reducing the risk of holes in the surface. We then gather the points on the remaining faces, setting their real values ($\psi$) to $1$, while setting the real values of other points to 00.
For 2D point cloud reconstruction, this process can be accelerated by ignoring faces that do not overlap significantly with the given point cloud. However, this approach is not applicable to 3D multi-view reconstruction.
@@
8.3 Step 2: Position Optimization
  • Line 3: We define the total number of optimization steps as $n_{0}$.
  • Line 5: We denote the number of nearest neighbors to store in the nearest neighbor cache for the query faces as $K$.
  • Lines 6-7: The optimization process runs for $n_{0}$ steps.
    • Extract points with a real value of $1$.
  • Lines 13-16: Using the current point configuration, we compute the minimum bounding balls ($B_{\mathbb{F}}$) for the query faces. For each bounding ball center, we find the nearest neighbor in the nearest neighbor cache $\mathbb{C}$ by calculating the distances to points in $\mathbb{C}$ and selecting the closest one. We then compute the signed distance $d(B_{\mathbb{F}},\mathbb{P})$ for the query faces and use it to determine the probability $\lambda_{min}(\mathbb{F})$.
@@ 8.3 Step 2: Position Optimization
  • Line 18: Based on the point positions, query faces, and their existence probabilities, we compute the loss $L$ to minimize. In this step, $L$ consists of the reconstruction loss $L_{recon}$ and the triangle quality loss $L_{qual}$ (for $d=3$). We calculate $L_{qual}$ as described in [37]. The total loss $L$ is defined as:
    $L=L_{recon}+\alpha_{qual}\cdot L_{qual},$ (17)
    where we set $\alpha_{qual}=10^{-3}$.
@@
8.4 Step 3: Point Reduction
In this step, we remove redundant points and faces using the Reinforce-Ball algorithm (Sec. 4.2) for $d=2$. For the formal definition of the algorithm, refer to Appendix 9.1.
8.5 Step 4: Real Value Optimization
@@
Next, we follow a similar optimization process to Step 1 (Appendix 8.2). Specifically, we optimize the face-wise probabilities using the same loss function and derive the point-wise real values from the optimized face-wise probabilities. In case of 3D multi-view reconstruction, we also conduct visibility test and remove faces that are not visible. After optimization, we optionally remove non-manifold structures, as described in Sec. 4.3, if desired by the user. If this is the last epoch, we return the post-processed mesh.
It is important to note that there may be faces that satisfy the tessellation function in Sec. 3.1 but are not included in the final mesh. This occurs because face-wise probabilities are optimized at this step, and non-manifold structures are removed based on face-wise existence.
8.6 Step 5: Subdivision
When $d=3$, we subdivide the mesh by inserting additional points on the current faces (Sec. 4.3, Appendix 10). However, for the 2D point cloud reconstruction task, we found that starting with a very dense grid eliminates the need for additional subdivisions.
9 Details about Reinforce-Ball algorithm
In the Reinforce-Ball algorithm, we select possible face combinations within the local neighborhood and then optimize the per-point probabilities ($\phi$) to minimize the number of points while preserving geometric details in 2D. Importantly, face combinations are determined locally, based on the Minimum-Ball condition, rather than globally. This local approach is the key to effectively removing redundant points and faces.
9.1 Algorithm
@@
  • Line 6: Gather the possibly existing faces ($\mathbb{F}$) based on the current point configuration and their real values. This function is the same as the one used in the Point Optimization step (Appendix 8.3).
@@
  • Line 13: For each batch, compute the loss as the sum of the reconstruction loss ($L_{recon}$) and the cardinality loss ($L_{card}$), as discussed in Sec. 4.2.

### v2

@@ 7.1 Algorithm
We formally describe the Minimum-Ball algorithm in Algorithm 1.
@@ 7.2 Minimum-Ball computation
$\scriptsize{B_{F}^{c}\big{|}_{d=3}=p_{1}+\frac{||d_{2}||^{2}(d_{1}\times d_{2})\times d_{1}+||d_{1}||^{2}(d_{2}\times d_{1})\times d_{2}}{2||d_{1}\times d_{2}||^{2}},}$ (7)
@@ 7.3 Sigmoid coefficient $\alpha_{min}$
As discussed in Appendix 8.2.1, the reconstruction process often starts from a fixed triangular (2D) or tetrahedral (3D) grid (Fig. 16). At the initial state, every face in the grid satisfies the Minimum-Ball condition (Appendix 8.2.1). Notably, every interior face in the grid shares a common signed distance $d_{common}>0$. Let us denote $x$ as the edge length of the grid, applicable for both 2D and 3D cases. Then, the common signed distance can be computed as follows:
@@ 7.4 Nearest neighbor caching
During the subsequent optimization steps, for a given face $F$, we compute the distance from $B_{F}^{c}$ to the cached $K$-nearest neighbors in $\mathbb{P}$ and select the nearest neighbor from the cache to compute the signed distance $d(B_{F},\mathbb{P})$. This mechanism is described in detail in Algorithm 2 and Appendix 8.2.2 in the context of point position optimization. In our experiments for 3D multi-view reconstruction, we set $n_{0}=2000$, $n_{1}=50$, and $K=10$.
Figure 14: Role of visibility gradient in geometric optimization. In this experiment, we optimize the translation vector of the object by comparing its rendered image and the ground truth image on the left. Since the differentiable renderer of DMesh [42] does not implement visibility gradient, while ours does, DMesh fails to find the correct translation vector.
8 Details about Reconstruction Process
In this section, we provide implementation details about our reconstruction process described in Sec. 4. Before delving into these details, we introduce the loss formulations for reconstruction problems.
8.1 Loss Formulation
Our final loss, $L$, is comprised of main reconstruction loss ($L_{recon}$) and two regularization terms: $L_{qual}$ and $L_{real}$.
$L=L_{recon}+\lambda_{qual}\cdot L_{qual}+\lambda_{real}\cdot L_{real}.$ (11)
We explain each of these terms below.
8.1.1 Reconstruction Loss ($L_{recon}$)
Reconstruction loss drives the reconstruction process by comparing our current probabilistic mesh and the given ground truth observations. For different observations, we need different loss functions as follows.
@@ 8.1.1 Reconstruction Loss ($L_{recon}$)
For rendering probabilistic faces, we interpret each face’s existence probability as its opacity, following [42]. To render large number of semi-transparent faces efficiently, we use the differentiable renderer of [42], but we found out that it does not implement visibility gradient that is necessary for optimizing geometric properties (Fig. 14). For the details about this visibility gradient, please refer to [19], which implemented the visibility gradient using anti-aliasing. Following their path, we enhanced the differentiable renderer of DMesh by implementing anti-aliasing in CUDA, which provides us visibility gradients.
In Fig. 15, we briefly illustrate how we implemented anti-aliasing in CUDA. Specifically, for each (triangular) face-pixel pair $(F,P)$, we project $F$ onto the image space and compute the overlapping area $A(F,P)$ between the projected triangle and the pixel (Fig. 15 right, blue area). Denoting the total area of the pixel as $A(P)$, the ratio of the overlapping area in the given pixel, $\rho(F,P)$, is computed as:
$\rho(F,P)=\frac{A(F,P)}{A(P)}.$ (12)
Figure 15: Implementation of anti-aliasing in our differentiable renderer. On the left, we show the process of anti-aliasing: 1) Find pixels that overlap with the given triangle, 2) Find the area that each pixel overlaps with the given triangle, and 3) Determine the color of each pixel based on the overlap area. On the right, we show details of the step 2.
$\alpha(F,P)=\alpha(F)\cdot\rho(F,P)\leq\alpha(F).$ (13)
In the right figure of Fig. 15, we illustrate the process of computing the overlapping area $A(F,P)$. The vertices of $F$ are projected onto the image plane and visited in counterclockwise order (e.g., A - B - C - A in the illustration). We then find the intersection points between the triangle edges and the pixel boundaries. These intersection points form the vertices of the (convex) overlapping polygon.
Using this enhanced differentiable renderer, we render multi-view images and compute the $L_{1}$ loss between the ground truth images as the reconstruction loss, $L_{recon}$.
8.1.2 Triangle Quality Loss ($L_{qual}$)
To improve the triangle quality of the final mesh, we adopt the triangle quality loss ($L_{qual}$) of DMesh [42]. Specifically, the loss is defined as follows:
$L_{qual}=\frac{1}{|\mathbb{F}|}\sum_{F\in\mathbb{F}}AR(F)\cdot\Lambda(F),$ (14)
where $\mathbb{F}$ is the set of every face combination we consider, $AR(\cdot)$ is a function that computes aspect ratio of the given face, and $\Lambda(\cdot)$ is the face probability function defined in Sec. 3.1.
8.1.3 Real Loss ($L_{real}$)
We minimize the sum of point-wise real values ($\psi)$, so that we can remove redundant faces as much as possible during optimization. The loss $L_{real}$ is simply defined as:
$L_{real}=\frac{1}{|\mathbb{P}|}\sum_{p\in\mathbb{P}}\Psi(p),$ (15)
where $\Psi(\cdot)$ is the function that returns the real value of the given point, as defined in Sec. 3.1.
8.2 Reconstruction Steps
Now we provide detailed explanations about each step in the reconstruction process.
8.2.1 Step 1: Initialization
We initialize our point features differently for two different scenarios: when sample point cloud is given or not.
Point Cloud Init.
When a point cloud sampled from the target shape is given, we can initialize our point features using the point cloud, so that the initial configuration already captures the overall structure of the target shape (Fig. 3). Specifically, for the given point cloud, we estimate the density of the point cloud by computing the distance to the nearest point for each point. Then, we down sample the point cloud using a voxel grid, of which size is defined as the point cloud density, to remove redundant points and holes. Finally, we follow the initialization scheme of DMesh [42] using the down sampled point cloud.
Regular Grid Init.
If we do not have any prior knowledge about the target shape, we first organize the points in a regular grid, ensuring that every face in the grid satisfies the Minimum-Ball condition (Definition 3.1), and initialize the point-wise real values ($\psi$) with additional features (e.g. colors). This regular grid guarantees that the faces observed in this step will also be observable in the subsequent step, where the Minimum-Ball algorithm determines face existence. For $d=2$, this condition is satisfied by forming every triangle in the grid as an equilateral triangle. For $d=3$, we use a body-centered cubic lattice. The grids are illustrated in Fig. 16.
With these fixed points and faces, we formulate the final loss by setting $\lambda_{qual}=0$ and $\lambda_{real}=10^{-4}$ in Eq. 11, and minimize it to determine which faces to include in the final mesh. After optimization, we collect points with real values larger than $0.01$ to ensure that as many faces as possible are available for the next optimization step, thereby reducing the risk of holes in the surface.
@@ 8.2.1 Step 1: Initialization
8.2.2 Step 2: Position Optimization
  • Line 3: We define the total number of optimization steps as $n_{0}$.
  • Line 5: We denote the number of nearest neighbors to store in the cache for the query faces as $K$.
  • Lines 6-7: The optimization process runs for $n_{0}$ steps.
    • Extract points with a real value of $1$.
@@ 8.2.2 Step 2: Position Optimization
  • Lines 13-16: Using the current point configuration, we compute the minimum bounding balls ($B_{\mathbb{F}}$) for the query faces. For each bounding ball center, we find the nearest neighbor in the nearest neighbor cache $\mathbb{C}$ by calculating the distances to points in $\mathbb{C}$ and selecting the closest one. We then compute the signed distance $d(B_{\mathbb{F}},\mathbb{P})$ for the query faces and use it to get the probability $\lambda_{min}(\mathbb{F})$.
@@
  • Line 18: Based on the point positions, query faces, and their existence probabilities, we compute the loss $L$ to minimize following Eq. 11.
@@
8.2.3 Step 3: Real Value Optimization
@@
Next, we follow a similar optimization process to Step 1 (Appendix 8.2.1). Additionally, if it was the multi-view reconstruction task, we remove invisible faces to remove redundant faces as much as possible. If this was the last epoch, we return the post-processed mesh.
8.2.4 Step 4: Subdivision
To reconstruct fine geometric details of the given shape, we subdivide the current mesh mainly by adding points with $\psi=1$ at the middle of edges that are adjacent to currently existing faces. In the inset, points $E,D,F$ are the newly inserted points. They form 4 sub-faces with $A,B,C$, and they all satisfy Definition 3.1. Therefore, we can guarantee that these sub-faces will exist at the start of next epoch. Note that this guarantee does not hold for WDT.
At the same time, it is also possible to insert additional points into faces that should not exist in the next epoch, effectively removing such faces at the start of the next epoch. For example, during the real value optimization step in the pipeline (Fig. 6, Appendix 8.2.3), invisible faces are removed for multi-view reconstruction task. After optimization, we may observe removed faces with all their points have a real value of $1.0$, creating a contradiction. This situation could arise due to ambiguities in the mesh definition, as illustrated in Fig. 17.
To eliminate these undesirable faces, additional points with a real value of 00 are inserted at their circumcenters. Consequently, after subdivision, several holes may appear on the surface because these additional points might also be inserted into faces that should exist (Fig. 6). However, most of these holes are resolved during subsequent optimization steps.
9 Experimental Details and Additional Results
In this section, we outline the experimental settings used for the results in Sec. 5 and present additional results to support our claims.
9.1 Dataset
Here, we provide details on the datasets described in Sec. 5.2.
9.1.1 Font
We used four font styles: Pacifico, Permanent-Marker, Playfair-Display, and Roboto.
9.1.2 Thingi10K
We manually selected 10 closed surfaces and 10 open surfaces from the Thingi10K dataset [52]. Specifically, we used the following models, denoted by their file IDs:
These models were chosen because they exhibit minimal self-occlusions, enabling dense observations from multi-view images. Additionally, we randomly selected 500 models and used for comparisons.
9.1.3 Objaverse
We manually selected 30 mesh models that exhibit diverse topology from Objaverse [10], which include both closed and open surfaces, and also have small scenes. Some of these models are rendered in Figs. 3, 6, 9, 20 and 19.
9.2 2D Point Cloud Reconstruction
9.2.1 Hyperparameters
DMesh++
  • Learning rate (real value, $\psi$): $0.3$
  • Learning rate (position): $0.001$
  • Number of epochs: 1
  • Number of optimization steps
    • Step 1 (Real value initialization): 100
    • Step 2 (Point position optimization): 500
9.2.2 Reconstruction of Complex Drawings
In Fig. 18, we present the reconstruction results for complex 2D drawings. As the figure illustrates, DMesh++ successfully reconstructs intricate 2D geometries from point clouds, even when the number of edges approaches nearly 1 million.
Figure 18: 2D point cloud reconstruction result for complex drawings. For each drawing, we report both the number of edges and the reconstruction time. For the Chinese drawing, we additionally render the “imaginary” part on the right to clearly illustrate its complexity.
9.3 3D Point Cloud Reconstruction
9.3.1 Hyperparameters
DMesh++
  • Initial Grid Edge Length: 3×3$\times$3 × input point cloud density
  • Learning rate (position): 0.001
  • Number of epochs: 1
  • Number of optimization steps
    • Step 2 (Point position optimization): 2000
    • Step 3 (Real value optimization): 0
9.4 3D Multi-View Reconstruction
Figure 19: Qualitative comparison of 3D point cloud reconstruction results for a closed surface (vase). For each image, we render the view-point normal on the right, and the diffuse image on the left. Among the baseline methods that reconstruct watertight mesh from point clouds, PoNQ [27] performs the best in reconstructing fine geometric details. While DMesh [42] fails at reconstructing such details due to the lack of mesh complexity, DMesh++ successfully recovers them and produce comparable result to PoNQ.
Figure 20: Qualitative comparison of 3D multi-view reconstruction results for a closed surface (sculpture). We render the input diffuse and depth images alongside the ground truth image. For each image, we render the view-point normal on the left, and the diffuse image on the right. We can observe that the reconstruction result of DMesh++ is as good as the other baseline methods that are optimized for closed surfaces.
9.4.1 Hyperparameters
Remeshing [33]
  • Image Batch Size: 8
  • Number of Optimization Steps: 1000
  • Learning Rate: 0.1
  • Edge Length Limits: [0.02, 0.15]
DMTet [38]
  • Image Batch Size: 8
  • Number of Optimization Steps: 5000
  • Learning Rate: 0.001
  • Grid Resolution: 128
FlexiCubes [39]
  • Image Batch Size: 8
  • Number of Optimization Steps: 2000
  • Number of Warm-up Steps: 1500
  • Learning Rate: 0.01
  • Grid Resolution: 80
  • Triangle Aspect Ratio Loss Weight: 0.01
GShell [23]
  • Image Batch Size: 8
  • Number of Optimization Steps: 5000
  • Number of Warm-up Steps: 4500
  • Learning Rate: 0.01
  • Grid Resolution: 80
  • Triangle Aspect Ratio Loss Weight: 0.0001
DMesh++ Settings
  • Initial Grid Edge Length: 0.05
  • Learning Rate (Real Value, $\psi$): 0.01
  • Learning Rate (Position): 0.001
  • Number of Epochs: 2
    • Image Res. / Batch Size at Epoch 1: (256, 256), 1
    • Image Res. / Batch Size at Epoch 2: (512, 512), 1
  • Number of Optimization Steps
    • Step 1 (Real Value Initialization): 1000
    • Step 2 (Point Position Optimization): 2000
    • Step 3 (Real Value Optimization): 1000
9.4.2 Limitations
Despite DMesh++’s success in reconstructing geometrically accurate meshes from multi-view images (of a synthetic object or scene), it currently cannot recover meshes from real-world images. This limitation stems from an inadequate rendering model — our current per-vertex color model is too simple to capture detailed geometry, and our algorithm assumes full knowledge of lighting conditions, which is not available in real-world scenarios.
Figure 21: 3D reconstruction from real-world images in DTU dataset [15]. The input images are shown on left, and the reconstructed mesh is shown on right.
To illustrate this, we applied our reconstruction algorithm to real-world images from the DTU dataset [15], as shown in Fig. 21. While our method approximates the real-world images (Fig. 21(b)), the extracted mesh exhibits numerous false floaters (Fig. 21(c)). We believe this suboptimal result is due to the lack of proper rendering models and regularizations, and addressing this issue by integrating DMesh++ with other reconstruction mechanisms is an exciting direction for future research, as discussed in Sec. 6.
10 Reinforce-Ball algorithm
Here we introduce an experimental algorithm that further enhances DMesh++’s capability. As discussed in Sec. 3.1, DMesh++ no longer uses the per-point weights found in DMesh [42]. In DMesh, optimizing per-point weights helps control mesh complexity: stronger regularization on these weights results in a simpler output mesh. Since DMesh++ lacks this mechanism, it cannot directly regulate mesh complexity during optimization. To address this limitation, we propose the Reinforce-Ball algorithm, which reduces unnecessary faces while preserving essential geometric details.
10.1 Local Minima of Weight Regularization
Before delving into the details of the Reinforce-Ball algorithm, we first highlight a limitation of DMesh’s per-point weight regularization. Specifically, while per-point weights are relevant for controlling mesh complexity, they alone cannot achieve adaptive resolution or produce a mesh that is both efficient and precise. Below, we explain the reasons in detail, assuming that per-point probabilities are optimized and that the Minimum-Ball condition is employed to compute face probabilities.
[…]
$0.2+(1.0-0.2)\cdot 1.0=1.0$ (17)
This dependency issue creates a local minimum that the previous formulation cannot overcome. This is why we propose the Reinforce-Ball algorithm.
10.2 Algorithm Overview
In the Reinforce-Ball algorithm, we define per-point existence probability and optimize it using stochastic optimization technique [48]. The overview of this algorithm is given in Fig. 23.
$L^{i}_{rl}=L^{i}_{recon}+\epsilon_{card}\cdot L^{i}_{card},$ (19)
where $\epsilon_{card}$ is a small tunable hyperparameter to adjust the weight of the cardinality loss. If we write the final loss for a set of sampled points $\mathbf{P}$ as $L_{rl}(\mathbf{P})$, we aim at minimizing the expected loss:
10.3 Formal Definition
@@
  • Line 6: Gather the possibly existing faces ($\mathbb{F}$) based on the current point configuration and their real values. This function is the same as the one used in the Point Optimization step (Appendix 8.2.2).
@@
    DMesh [42] (0) 1.97 2506 2245 30.39
    DMesh ($10^{-4}$) 2.68 666 693 153.10
    DMesh ($10^{-3}$) 12.48 456 488 152.37
    DMesh++ (0) 1.82 2862 2793 11.33
    DMesh++ ($10^{-6}$) 1.86 1386 1394 278.88
    DMesh++ ($10^{-5}$) 2.77 149 152 200.05
@@
  • Line 13: For each batch, compute the loss as the sum of the reconstruction loss ($L_{recon}$) and the cardinality loss ($L_{card}$), as discussed in Appendix 10.2.